Opened 9 years ago

Closed 3 years ago

#343 closed flight processing (fixed)

GB08/19, 154a/2010, London day

Reported by: jb362 Owned by: knpa
Priority: alpha 5 Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: iopa@… Other processors: iopa

Description (last modified by emca)

Data location: users/rsg/airborne/workspace/GB08_19-2010_154a_London_day

Data arrived from ARSF via network transfer on 08/06/2010.

Scientific objectives:

  1. explore the use of airborne remote sensing (hyperspectral imaging, thermal infrared measurement and LiDAR) for paramterising models of the urban surface energy balance
  2. to determine the degree of agreement between the output of such models and traditional in situ energy balance measures (which provide excellent accuracy but poor or no spatial detail)
  3. to assess the thermal anisotropy of the three dimensional urban surface structure as it relates to modelling urban energy balances and observing the urban radiation balance
  4. use the same data to explore patterns of urban heat island intensity that directly impact human comfort/health

Priority: a5

PI: M. J. Wooster

Sensors:

  • Eagle (13/08/2010)
  • Hawk (20/12/2010)
  • Leica LIDAR (13/08/2010)
    • NB. Phil would like a screenshot of the joined LAS lines from this one for internal use - please ask Mike about it when you get to a LIDAR delivery.
  • Photography (13/08/2010)

Attachments (1)

1542010.sync (106 bytes) - added by iopa 8 years ago.
Eagle & Hawk SCT offsets

Download all attachments as: .zip

Change History (46)

comment:1 Changed 9 years ago by jb362

This directory needs to be backed up in the repository

comment:2 Changed 9 years ago by jb362

IPAS navigation for this project has been processed.

comment:3 Changed 9 years ago by chrfi

  • Owner set to chrfi
  • Status changed from new to assigned

processing lidar

comment:4 Changed 9 years ago by ella

Applanix navigation processing now done. Files saved in ARSF wirkspace on Samba.

comment:5 Changed 9 years ago by iopa

  • Cc iopa@… added
  • Other processors set to iopa

Hyperspectral processing picked up by me.

comment:6 follow-up: Changed 9 years ago by iopa

  1. Only 5 lines recorded in log sheet vs. 6 lines recorded in eagle/hawk directories.
  1. First eagle line shows different times in the .hdr compared to the sbet file. -> .hdr displays only 14 seconds of data

Will try to discard line 1 which has almost no raw data recorded.

comment:7 in reply to: ↑ 6 Changed 9 years ago by iopa

-> the above is the case for both eagle and hawk data, so only lines 2-6 in eagle/hawk (all lines in logsheet) will be processed

comment:8 Changed 9 years ago by mggr

  • Description modified (diff)

Phil would like a screenshot of the joined LAS lines from this one for internal use - please ask Mike about it when you get to a LIDAR delivery.

comment:9 Changed 9 years ago by iopa

The PI, M. Wooster, has requested us to CC this person ("Kotthaus,Simone" <simone.kotthaus@…>) regarding any developments in this project.

comment:10 Changed 9 years ago by chrfi

lidar proccesing complete, awaiting delivery check

comment:11 Changed 9 years ago by iopa

The PI was emailed and replied saying that they are willing to wait for vectors to verify the data.

comment:12 Changed 9 years ago by emca

Delivery checked the LIDAR data at
~airborne/workspace/GB08_19-2010_154a_London_day/delivery/20100709/

Screenshot of intensity needs to be added to the delivery. Otherwise fine.

As per above comment PI would like to wait for vectors before sending out the data.

comment:13 Changed 9 years ago by emca

Phil no longer requires a screen shot of the LIDAR data.
Gary has requested some RCD images over Westminster/London Eye to be sent to him on disk.

comment:14 Changed 9 years ago by emca

The data for this flight does not cover the area requested by Gary, however flight 154b (ticket #344) does. As there are no RCD images for that flight (it was a night flight) will send a LIDAR intensity image covering the requested area instead.

comment:15 Changed 9 years ago by knpa

  • Owner changed from chrfi to knpa

Taking over responsibility for LIDAR as chrfi is leaving.

comment:16 Changed 9 years ago by emca

  • Description modified (diff)

Photography data processed and awaiting delivery check at
~airborne/workspace/GB08_19-2010_154a_London_day/delivery/20100729

comment:17 Changed 9 years ago by emca

Removed photographs 1-14 as they did not correspond to the area of interest.

comment:18 Changed 9 years ago by jb362

Beggining camera delivery check

comment:19 Changed 9 years ago by jb362

Camera delivery check complete and everything looks good except according to the wiki the .kml file should be renamed from GB08_19-2010154-Photographs.kml to GB08_19-2010_154a_London_day.kml

comment:20 Changed 9 years ago by knpa

  • Owner changed from knpa to emca

emca is taking over lidar

comment:21 Changed 9 years ago by iopa

Hyperspectral delivery ready to be checked.

comment:22 Changed 9 years ago by anch

I have checked the hyperspectral and everything looks fine.

I haven't checked the Hawk data due to the band issue.

comment:23 Changed 9 years ago by iopa

A disk was dispatched today to Prof. M. Wooster containing the processed LIDAR, Eagle and Photography data. Hawk was not delivered as of yet due to known calibration issue.

comment:24 Changed 9 years ago by iopa

Date of dispatch : 13th August 2010

The disk also included data from the project GB07/12 Debden Julian day 99, data included LIDAR and Eagle.

PI will shortly be emailed of the dispatch.

comment:25 Changed 9 years ago by emca

  • Description modified (diff)

comment:26 Changed 8 years ago by emca

Delivery checked Hawk.
Flight letter needs to be added to filenames + readme

Otherwise good to deliver.

comment:27 Changed 8 years ago by emca

  • Description modified (diff)

Hawk data sent 08/12/2010

Changed 8 years ago by iopa

Eagle & Hawk SCT offsets

comment:28 Changed 8 years ago by iopa

Hawk data is being re-processed for new calibration wavelength set.

comment:29 Changed 8 years ago by iopa

New Hawk data has been re-processed and added in the delivery 20101216/.

comment:30 Changed 8 years ago by iopa

Hawk data has been checked and is ready to deliver.

comment:31 Changed 8 years ago by emca

  • Description modified (diff)

Hawk data re-delivered 20/12/2010

comment:32 Changed 8 years ago by knpa

  • Owner changed from emca to knpa

Raw rcd image and log files for 154b were in here, have moved to correct project.

comment:33 Changed 8 years ago by knpa

  • Component changed from Processing: general to Archiving

R-synced whole project.

Beginning archiving.

comment:34 Changed 8 years ago by knpa

Tabi data present for this project.

comment:35 Changed 8 years ago by knpa

Some raw FW lidar present for this project.

comment:36 Changed 8 years ago by knpa

Issue with delivery 20100713, seems to have included un-calibrated hawk data and Readme for wrong project. iopa delivered this so will ask him when he's back.

comment:37 Changed 8 years ago by knpa

iopa has e-mailed PI to ask if they have correct Readme and telling them to disregard any hawk files that were sent with the earlier eagle delivery.

Will hold-off on archiving this project until we get a reply and know if we need to regenerate the Readme.

comment:38 Changed 7 years ago by knpa

Re-picking up the archiving on this now that 2011 is done. There is a new hawk delivery with presumably corrected Hawk data. The ReadMe for the original HS delivery is still missing.

comment:39 Changed 7 years ago by knpa

  • Milestone set to 2010 data processing completion

comment:40 Changed 7 years ago by knpa

Correction - the updated eagle delivery was in the workspace. Have replaced the one in the repository.

comment:41 Changed 7 years ago by knpa

Made available to NEODC.

comment:42 Changed 7 years ago by knpa

Ingested by NEODC.

Moved to archive disk.

comment:43 Changed 6 years ago by knpa

Removed workspace version (original processing).

Leaving ticket open until reprocessing is archived.

comment:44 Changed 5 years ago by knpa

Archiving reprocessed HS.

There were two of these deliveries. Moved the most recent one to delivery/ and moved GB08_19-154a-hyperspectral-20120419/ to processing/deliveries/old_deliveries/

comment:45 Changed 3 years ago by dac

  • Resolution set to fixed
  • Status changed from assigned to closed
Note: See TracTickets for help on using tickets.